Preliminary Thoughts on Rush Improvement

Random Hall MRNSA Task Force

September 4, 2002

1 Goals

1. Nothing else scheduled during Rush (esp. tests)

2. Make the Adjustment Lottery mandatory

3. Encourage dorm Rush

4. Make a dorm referral (Flush) book

5. Improve handling of RBA

6. Improve website and summer mailings

2 MRNSA timeline

1. DormCon meeting

2. Meetings in other dorms

3. All-living group meeting

4. Talk with UA and GSC

5. Talk with Tony Gray (RLA)

6. Write documents for admins.

7. Talk to larryben and Redwine

8. Talk to Corporation

3 What Went Right

• Midway — new format was an improvement

• Dead week

• Good Orientation leaders (especially from frats)

• Always something to do (Dorm activities)

• Got to meet people from all over campus

• Dinner w/profs

• Moving frosh into temp rooms (get to meet them)

• Early housing info for temps makes parents happy

4 Problems in the Administration

• Unclear communication to frosh over the summer • Lack of firm, clear deadlines

• Advanced standing exams during Rush (and at 0800)

• Only allowed to rank 3 choices in Adjusment Lottery • Poor 2006 website (construction and maintenance)

• Adjustment Lottery results not pre-checked by RACs • Parents were around and influential during Rush • Lack of information about RBA over summer

• Lack of communication to dorm Rush chairs

5 Possible Solutions

• Coordinate all mailings through one office

• Have advanced standing exams done before rush • Reformulate adjustment lottery for many moves

• Employ a student (or a better one) to handle website • Move Parents’ Weekend to end of Rush

• Let people escape RBA (by moving or dropping)

• Send all information going to frosh to rush chairs